# disposition

### v2

@@ 6.3 Fisher Transformations
for a generic (population or estimated) dependence measure $\delta\in\{{Q},{C},\phi,\widehat{{Q}}_{n},\widehat{{C}}_{n},\widehat{\phi}_{n}\}$. The $\mathrm{arctanh}(\cdot)$ in (10) bijectively maps the interval $(-1,1)$ to $\mathbb{R}$ and hence provides a better approximation when the dependence measure is close to the boundary.
@@ 7 Case Study on Drug Use
We use data from the 2018 wave of the Monitoring the Future (MTF) study on drug use among the contemporary American youth (Miech et al.,, 2019, DS1 Core Data). [^3] We use data from 2018 because subsequent waves have sizeably fewer observations. We are interested in interdependence between consumption of different drugs and consider the events of ever having consumed certain drugs. The sample size is $14{,}502$, but for certain drugs there is a large number of non-responses. [^4] The dataset is an aggregation of six different questionnaires students get randomly assigned to. Some drugs do not appear in all six questionnaires. The different drugs, the abbreviations we use for them in subsequent plots, their marginal relative consumption frequencies and the respective (univariate) sample sizes are listed in Table 4.
@@
alcohol ALC $0.60$ $13{,}219$
marijuana MAR $0.44$ $13{,}395$
cigarettes CIG $0.24$ $13{,}680$
amphetamines AMP $0.09$ $13{,}458$
tranquilizers TRQ $0.07$ $13{,}408$
narcotics NAR $0.06$ $13{,}305$
LSD LSD $0.05$ $11{,}210$
psychedelics PSY $0.04$ $11{,}202$
sedatives SED $0.04$ $13{,}476$
MDMA MDM $0.04$   $6{,}736$
cocaine COK $0.04$ $13{,}361$
crack CRA $0.01$ $12{,}920$
methamphetamines MET $0.01$   $4{,}446$
heroin HER $0.01$ $13{,}444$
@@
Figure 4 shows contour plots of estimated correlation matrices between all pairs of those events for Cole’s ${C}$, Yule’s ${Q}$ and the phi coefficient. All three measures indicate a positive relationship between the consumption of all the drugs considered. However, most values of the phi coefficient are very close to 0, which would usually be interpreted as evidence for very weak dependence, while the proper dependence measures tend to have very high values, indicating strong dependence. Thus, the dependence is structurally underrated by the phi coefficient due to its severe attainability issues. If the phi coefficient or one of the equally popular contingency coefficients was used in such a context, it could thus lead to possibly severe misperceptions with respect to the nature of polydrug use.
@@
While the two proper correlation matrices all indicate a very strong dependence overall, there are nevertheless clearly visible differences between them. Those differences can be very well understood by looking again at Figure 2. For example, Yule’s ${Q}$ takes considerably higher values than Cole’s ${C}$ for all drugs except for the three softest. This is due to the marginal event probabilities being rather small and similar, corresponding to a situation as in the lower left corner of the corresponding panels of Figure 2. Thus, this data example also illustrates that even though the proper dependence measures are usually close to each other (see again Figure 2), there may be substantial differences between them.
To get an idea of sampling uncertainty, we first focus on the dependence between marijuana and all other drugs (the second row or column of the correlation matrices above) and plot $\widehat{{C}}_{n}$, $\widehat{\phi}_{n}$ and $\widehat{{Q}}_{n}$ and 90% confidence intervals based on the Fisher transformation in Figure 5. Even though the consumption events for the harder drugs are rather rare events (that is, the marginal and even more so the joint consumption probabilities can be quite small), the confidence intervals are quite narrow due to the large sample size. As expected, the intervals for ${C}$ are slightly wider than the intervals for ${Q}$ due to their more complicated and possibly conservative construction method. Notice that the narrow confidence intervals for $\phi$ are not an advantage of this measure but are caused by the suboptimal normalization of $\phi$ that forces $\phi$ to be much smaller in absolute value due to its non-attainability.
@@ 8 Conclusion
This paper deals with dependence measures for events or binary random variables. We introduce dependence concepts for events and the notion of a proper dependence measure. We then discuss theoretical properties of, statistical inference for, and relations between different measures. Important recommendations for statistical practice arise: The most widely-used measures, the phi coefficient and the closely related contingency coefficients, should be avoided for this purpose. They lack the crucial property of attainability and can hence severely understate strength of dependence, leading to possibly misleading conclusions. Instead, proper dependence measures should be used. The largely unknown Cole coefficient ${C}$ is a very natural and nicely interpretable representative of this class. Even though its statistical inference is challenging due to the case distinction in its definition, we construct tests and confidence intervals based on the asymptotic distribution that we derive. Yule’s ${Q}$ is an attractive alternative, avoiding the case distinction in the normalization, and making possible the use of classical inferential procedures. We further discuss a generalized Yule’s ${Q}$, the also closely related and popular odds ratio and tetrachoric correlation and establish their propriety.
@@ A.1 Contingency Coefficients
Several contingency coefficients have been defined based on Pearson’s Mean Square Contingency coefficient ${MSC}(X,Y)$ from (4) (see Liebetrau, (1983) for an overview), including Cramérs ${V}$, Tschuprow’s ${T}$, and Pearson’s contingency coefficient ${PC}$
@@
We write ${V}(A,B):={V}(\mathds{1}_{A},\mathds{1}_{B})$ and analogously ${T}(A,B)$ and ${PC}(A,B)$. Here, Cramérs ${V}$ and Tschuprow’s ${T}$ coincide, ${V}(A,B)={T}(A,B)=\sqrt{{MSC}(A,B)}$. Thus, the contingency coefficients are just functions of the phi coefficient.
@@
Thus, in the binary case, Cramérs ${V}$ and Tschuprow’s ${T}$ just report the same information on strength of dependence as the phi coefficient, but remove the sign, that is, the information on direction of dependence. They of course inherit the attainability problems of $\phi(A,B)$ and are thus no proper dependence measures (not even in terms of a modified set of axioms adopted to measures of strength of dependence mapping to $[0,1]$). Pearson’s contingency coefficient ${PC}(A,B)$ exhibits even more shortcomings, for example its maximum value is $\frac{1}{\sqrt{2}}$ instead of 1.
@@ A.2 Odds Ratio
The origin of the odds ratio seems to be unknown, but Cornfield, (1951) has played an important role in popularizing the measure in epidemiology (Greenhouse,, 1982). It is very popular in medicine. Indeed, it surmounts all the classical dependence measures discussed here in terms of Google $(75{,}200{,}000)$ and Google Scholar $(1{,}440{,}000)$ citations (30th October 2024, compare Table 2), presumably due to the sheer size of the medical literature. The quotient between a probability (here: $\mathbb{P}(A|B)$ or $\mathbb{P}(B|A)$) and its counterprobability is called the odds. Dividing the odds for $A$ conditional on $B$ and the odds for $A$ conditional on $\overline{B}$ by each other yields the odds ratio,
@@
    Independence: ${OR}(A,B)=1$ if and only if $A$ and $B$ are independent.
@@
    Symmetry: ${OR}(A,B)={OR}(B,A)$ and ${OR}(A,\overline{B})=\frac{1}{{OR}(A,B)}$.
A.3 Invariance to the Marginals of ${Q}$, ${Q}_{g}$ and the Odds Ratio
${Q}$, ${OR}$ and ${Q}_{g}$ have an interesting property, which is discussed already in Yule, (1912) and later in Edwards, (1963), Plackett, (1965) and Mosteller, (1968) for empirical contingency tables containing absolute frequencies: When the columns and rows of a $2\times 2$ contingency table (see Table 3 again) are multiplied by constants $c_{1},c_{2}$ and $r_{1},r_{2}$ and then the numbers are normalized by an appropriate normalising constant $\nu$ so that the joint probabilities sum up to 1 again, those measures do not change. Considering the odds ratio, it is easy to see that those constants cancel out:
@@ A.3 Invariance to the Marginals of ${Q}$, ${Q}_{g}$ and the Odds Ratio
By Lemma 5.6, this directly carries over to Yule’s ${Q}$. This property is seen as a desirable property by the papers cited above and often described as invariance of those measures to the marginal event probabilities because the measures do not change when the marginal event probabilities change and the respective joint and conditional probabilities change accordingly. On the one hand, this is indeed a very nice property as it allows to really disentangle the marginal distributions of the events and their dependence. On the other hand, this property is incompatible with the property that after distinguishing positive and negative dependence the dependence measure is proportional to ${Cov}(A,B)$ and a linear function of $\mathbb{P}(A\cap B)$. This is a property that only Cole’s ${C}$ possesses, whereas the normalization of ${Q}$ changes with the joint probabilities. We do not take a strong stance regarding those properties. Both measures fulfill the more fundamental properties leading to propriety and both are very natural dependence measures.
@@ A.4 Tetrachoric Correlation
For interior cases $(\max(\mathbb{P}(A)+\mathbb{P}(B)-1,0)<\mathbb{P}(A\cap B)<\min(\mathbb{P}(A),\mathbb{P}(B)))$, tetrachoric correlation ${TC}(A,B)$ is implicitly defined via the equation
[…]
where $f_{\rho}$ denotes the density of a bivariate standard normal distribution with zero means, unit variances and correlation coefficient $\rho$. Otherwise, we have ${TC}(A,B)=-1$ for $\mathbb{P}(A\cap B)=\max(\mathbb{P}(A)+\mathbb{P}(B)-1,0)$ and ${TC}(A,B)=1$ for $\mathbb{P}(A,B)=\min(\mathbb{P}(A),\mathbb{P}(B))$.
@@
${TC}(A,B)$ is proper.
@@ Appendix B Inference Using the Fisher Transformation
for a generic dependence measure $\delta\in\{{Q},{C},\phi,\widehat{{Q}}_{n},\widehat{{C}}_{n},\widehat{\phi}_{n}\}$. This transformation maps the bounded dependence measures to an unbounded scale while leaving values in the center (approximately the interval $[-0.5,0.5]$) almost unchanged. On this unbounded scale, asymptotic normality is usually a much better finite sample approximation, especially when the dependence measure is close to the boundary. We therefore apply the Fisher transformation to construct more reliable tests and confidence intervals for ${Q}$ (as brought up in Bonett and Price, (2007)), ${C}$ and $\phi$. Even though the limiting distribution is not Gaussian for ${C}$, the same problem applies here (unless ${C}=0$).
@@
Starting with the limiting distributions of the Fisher transformation as given below in Propositions B.1B.3, we obtain non-normal limiting distributions for $\widehat{{Q}}_{n}$, $\widehat{{C}}_{n}$ and $\widehat{\phi}_{n}$ by applying the inverse transformation $Z^{-1}(z)=\mathrm{tanh}(z)=(e^{2z}-1)/(e^{2z}+1)$ of (13). These usually lead to much better approximations of the sampling distributions close to the boundaries as illustrated in Figure 6. For the true ${Q}=0.95$ with $p=q=0.5$, we plot the standard (red) and Fisher transformed (blue) limiting distributions, their implied $90\%$ confidence intervals in dashed lines, and a histogram depicting $100{,}000$ estimates $\widehat{{Q}}_{n}$ for a sample size of $n=100$ from simulated data. We see that the finite sample histogram is much better approximated by the blue Fisher transformed limiting distribution.
@@
Figure 6: Histogram of $\widehat{{Q}}_{n}$ from $100{,}000$ simulated samples of size $n=100$ with a true value of ${Q}=0.95$, marginal probabilities $p=q=0.5$ and the standard asymptotically normal and the (backtransformed) Fisher transformed limiting distributions together with corresponding $90\%$ confidence intervals in dashed lines.
@@
Consequently, the Fisher transformation leads to more reliable tests and confidence intervals close to the boundaries and almost identical results away from them. To execute a test for $\mathcal{H}_{0}:\delta=d$ one just needs to Fisher transform the hypothesized value and test $\mathcal{H}_{0}:Z(\delta)=Z(d)$ using the respective asymptotic distribution of the Fisher transformation (again using Bonferroni corrections as described above in the case of ${C}$).
@@
We construct confidence intervals as follows. For ${Q}$ and $\phi$, we obtain a confidence interval for the Fisher transform based on its asymptotic normality and then apply the inverse Fisher transformation to these bounds. For ${C}$, we modify the algorithm described in Section 6.2 by using (inverted) tests based on the asymptotic distribution of the Fisher transforms from Proposition B.2 instead of Proposition 6.5. We illustrate the improved performance of the confidence intervals based on the Fisher transform compared to the ones based on the usual limiting distributions in simulations in Appendix D. The importance of the Fisher transformation is also illustrated when comparing the confidence intervals based on it to standard confidence intervals in the introductory example, where they differ substantially as the sampling distribution is close to the upper bound 1, see Table 6 in the Appendix, and in the case study in Section 7, where they differ in those cases and are very close to each other otherwise, see Figures 13 and 15 in the Appendix.
@@
We continue by formally providing the limiting distributions of the Fisher transformations of ${Q}$, ${C}$ and $\phi$ introduced in Section 6.3, which can be obtained from the asymptotic distributions from Subsection 6.1 together with the delta method. For Yule’s ${Q}$, we instead derive the limit distribution directly as in this case the transformation is a simpler expression than Yule’s ${Q}$ itself, namely $Z({Q})=\frac{1}{2}\log({OR})$ with the plug-in estimator
@@
Furthermore, its first row simplifies to $-\frac{1-\mathsf{q}}{2(\mathsf{p}-\mathsf{r})(1-\mathsf{p}-\mathsf{q}+\mathsf{r})}$, which is nonzero by Assumption 6.1. Hence, by the delta method (van der Vaart,, 1998, Theorem 3.1) and Lemma E.2 the claim follows. ∎
@@
[…]
@@
    • if $p=1-q$, then $\sqrt{n}\big{(}Z(\widehat{{C}}_{n})-Z({C})\big{)}\stackrel{{\scriptstyle d}}{{\to}}\gamma_{C}\Lambda^{-}\begin{pmatrix}V_{4}\\ V_{3}-\mathds{1}\{V_{1}>-V_{2}\}\big{[}V_{1}+V_{2}\big{]}\end{pmatrix}$.

### v3

@@ 6.3 Fisher Transformations
for a generic (population or estimated) dependence measure $\delta\in\{{Q},{C},\phi,\widehat{{Q}}_{n},\widehat{{C}}_{n},\widehat{\phi}_{n}\}$. The $\mathrm{arctanh}(\cdot)$ in (10) bijectively maps the interval (1,1)(-1,1) to $\mathbb{R}$ and hence provides a better approximation when the dependence measure is close to the boundary.
@@ 7 Case Study on Drug Use
We use data from the 2018 wave of the Monitoring the Future (MTF) study on drug use among the contemporary American youth (Miech et al.,, 2019, DS1 Core Data). [^3] We use data from 2018 because subsequent waves have sizeably fewer observations. We are interested in interdependence between consumption of different drugs and consider the events of ever having consumed certain drugs. The sample size is 14,50214{,}502, but for certain drugs there is a large number of non-responses. [^4] The dataset is an aggregation of six different questionnaires students get randomly assigned to. Some drugs do not appear in all six questionnaires. The different drugs, the abbreviations we use for them in subsequent plots, their marginal relative consumption frequencies and the respective (univariate) sample sizes are listed in Table 4.
@@
alcohol ALC $0.60$ 13,21913{,}219
marijuana MAR $0.44$ 13,39513{,}395
cigarettes CIG $0.24$ 13,68013{,}680
amphetamines AMP $0.09$ 13,45813{,}458
tranquilizers TRQ $0.07$ 13,40813{,}408
narcotics NAR $0.06$ 13,30513{,}305
LSD LSD $0.05$ 11,21011{,}210
psychedelics PSY $0.04$ 11,20211{,}202
sedatives SED $0.04$ 13,47613{,}476
MDMA MDM $0.04$   6,7366{,}736
cocaine COK $0.04$ 13,36113{,}361
crack CRA $0.01$ 12,92012{,}920
methamphetamines MET $0.01$   4,4464{,}446
heroin HER $0.01$ 13,44413{,}444
@@
Figure 4 shows contour plots of estimated correlation matrices between all pairs of those events for Cole’s C{C}, Yule’s Q{Q} and the phi coefficient. All three measures indicate a positive relationship between the consumption of all the drugs considered. However, most values of the phi coefficient are very close to 0, which would usually be interpreted as evidence for very weak dependence, while the proper dependence measures tend to have very high values, indicating strong dependence. Thus, the dependence is structurally underrated by the phi coefficient due to its severe attainability issues. If the phi coefficient or one of the equally popular contingency coefficients was used in such a context, it could thus lead to possibly severe misperceptions with respect to the nature of polydrug use.
@@
While the two proper correlation matrices all indicate a very strong dependence overall, there are nevertheless clearly visible differences between them. Those differences can be very well understood by looking again at Figure 2. For example, Yule’s Q{Q} takes considerably higher values than Cole’s C{C} for all drugs except for the three softest. This is due to the marginal event probabilities being rather small and similar, corresponding to a situation as in the lower left corner of the corresponding panels of Figure 2. Thus, this data example also illustrates that even though the proper dependence measures are usually close to each other (see again Figure 2), there may be substantial differences between them.
To get an idea of sampling uncertainty, we first focus on the dependence between marijuana and all other drugs (the second row or column of the correlation matrices above) and plot $\widehat{{C}}_{n}$, $\widehat{\phi}_{n}$ and $\widehat{{Q}}_{n}$ and 90% confidence intervals based on the Fisher transformation in Figure 5. Even though the consumption events for the harder drugs are rather rare events (that is, the marginal and even more so the joint consumption probabilities can be quite small), the confidence intervals are quite narrow due to the large sample size. As expected, the intervals for C{C} are slightly wider than the intervals for Q{Q} due to their more complicated and possibly conservative construction method. Notice that the narrow confidence intervals for $\phi$ are not an advantage of this measure but are caused by the suboptimal normalization of $\phi$ that forces $\phi$ to be much smaller in absolute value due to its non-attainability.
@@ 8 Conclusion
This paper deals with dependence measures for events or binary random variables. We introduce dependence concepts for events and the notion of a proper dependence measure. We then discuss theoretical properties of, statistical inference for, and relations between different measures. Important recommendations for statistical practice arise: The most widely-used measures, the phi coefficient and the closely related contingency coefficients, should be avoided for this purpose. They lack the crucial property of attainability and can hence severely understate strength of dependence, leading to possibly misleading conclusions. Instead, proper dependence measures should be used. The largely unknown Cole coefficient C{C} is a very natural and nicely interpretable representative of this class. Even though its statistical inference is challenging due to the case distinction in its definition, we construct tests and confidence intervals based on the asymptotic distribution that we derive. Yule’s Q{Q} is an attractive alternative, avoiding the case distinction in the normalization, and making possible the use of classical inferential procedures. We further discuss a generalized Yule’s Q{Q}, the also closely related and popular odds ratio and tetrachoric correlation and establish their propriety.
@@ A.1 Contingency Coefficients
Several contingency coefficients have been defined based on Pearson’s Mean Square Contingency coefficient MSC(X,Y){MSC}(X,Y) from (4) (see Liebetrau, (1983) for an overview), including Cramérs V{V}, Tschuprow’s T{T}, and Pearson’s contingency coefficient PC{PC}
@@
We write ${V}(A,B):={V}(\mathds{1}_{A},\mathds{1}_{B})$ and analogously T(A,B){T}(A,B) and PC(A,B){PC}(A,B). Here, Cramérs V{V} and Tschuprow’s T{T} coincide, ${V}(A,B)={T}(A,B)=\sqrt{{MSC}(A,B)}$. Thus, the contingency coefficients are just functions of the phi coefficient.
@@
Thus, in the binary case, Cramérs V{V} and Tschuprow’s T{T} just report the same information on strength of dependence as the phi coefficient, but remove the sign, that is, the information on direction of dependence. They of course inherit the attainability problems of $\phi(A,B)$ and are thus no proper dependence measures (not even in terms of a modified set of axioms adopted to measures of strength of dependence mapping to $[0,1]$). Pearson’s contingency coefficient PC(A,B){PC}(A,B) exhibits even more shortcomings, for example its maximum value is $\frac{1}{\sqrt{2}}$ instead of 1.
@@ A.2 Odds Ratio
The origin of the odds ratio seems to be unknown, but Cornfield, (1951) has played an important role in popularizing the measure in epidemiology (Greenhouse,, 1982). It is very popular in medicine. Indeed, it surmounts all the classical dependence measures discussed here in terms of Google (14,200,000)(14{,}200{,}000) and Google Scholar (1,590,000)(1{,}590{,}000) citations (7th November 2025, compare Table 2), presumably due to the sheer size of the medical literature. The quotient between a probability (here: $\mathbb{P}(A|B)$ or $\mathbb{P}(B|A)$) and its counterprobability is called the odds. Dividing the odds for $A$ conditional on $B$ and the odds for $A$ conditional on $\overline{B}$ by each other yields the odds ratio,
@@
    Independence: OR(A,B)=1{OR}(A,B)=1 if and only if $A$ and $B$ are independent.
@@
    Symmetry: OR(A,B)=OR(B,A){OR}(A,B)={OR}(B,A) and ${OR}(A,\overline{B})=\frac{1}{{OR}(A,B)}$.
A.3 Invariance to the Marginals of Q{Q}, ${Q}_{g}$ and the Odds Ratio
Q{Q}, OR{OR} and ${Q}_{g}$ have an interesting property, which is discussed already in Yule, (1912) and later in Edwards, (1963), Plackett, (1965) and Mosteller, (1968) for empirical contingency tables containing absolute frequencies and plays a central role in an approach to copula modeling and measuring dependence suggested in Geenens, (2020) and Geenens, (2023): When the columns and rows of a $2\times 2$ contingency table (see Table 3 again) are multiplied by constants $c_{1},c_{2}$ and $r_{1},r_{2}$ and then the numbers are normalized by an appropriate normalising constant $\nu$ so that the joint probabilities sum up to 1 again, those measures do not change. Considering the odds ratio, it is easy to see that those constants cancel out:
@@ A.3 Invariance to the Marginals of Q{Q}, ${Q}_{g}$ and the Odds Ratio
By Lemma 5.6, this directly carries over to Yule’s Q{Q}. This property is seen as a desirable property by the papers cited above and often described as invariance of those measures to the marginal event probabilities because the measures do not change when the marginal event probabilities change and the respective joint and conditional probabilities change accordingly. On the one hand, this is indeed a very nice property as it allows to really disentangle the marginal distributions of the events and their dependence. On the other hand, this property is incompatible with the property that after distinguishing positive and negative dependence the dependence measure is proportional to Cov(A,B){Cov}(A,B) and a linear function of $\mathbb{P}(A\cap B)$. This is a property that only Cole’s C{C} possesses, whereas the normalization of Q{Q} changes with the joint probabilities. We do not take a strong stance regarding those properties. Both measures fulfill the more fundamental properties leading to propriety and both are very natural dependence measures.
@@ A.4 Tetrachoric Correlation
For interior cases $(\max(\mathbb{P}(A)+\mathbb{P}(B)-1,0)<\mathbb{P}(A\cap B)<\min(\mathbb{P}(A),\mathbb{P}(B)))$, tetrachoric correlation TC(A,B){TC}(A,B) is implicitly defined via the equation
$\mathbb{P}(A\cap B)=\int_{-\infty}^{\Phi^{-1}(P(A))}\int_{-\infty}^{\Phi^{-1}(P(B))}f_{{TC}(A,B)}(x,y)dxdy,$
where $f_{\rho}$ denotes the density of a bivariate normal distribution with zero means, unit variances and correlation coefficient $\rho$. Otherwise, we have TC(A,B)=1{TC}(A,B)=-1 for $\mathbb{P}(A\cap B)=\max(\mathbb{P}(A)+\mathbb{P}(B)-1,0)$ and TC(A,B)=1{TC}(A,B)=1 for $\mathbb{P}(A,B)=\min(\mathbb{P}(A),\mathbb{P}(B))$.
@@
TC(A,B){TC}(A,B) is proper.
A.5 Further Measures
In this subsection, we collect further measures that have originally not been designed for the binary case. Still, they can be used and their popularity in the contemporaneous statistical literature justifies a separate treatment.
The first of such measures is distance correlation (Székely et al.,, 2007). Its numerator, the distance covariance, is defined as
$\mathcal{V}(X,Y):=\frac{1}{c_{p}c_{q}}\int_{\mathbb{R}^{p+q}}\frac{|f_{X,Y}(t,s)-f_{X}(t)f_{Y}(s)|^{2}}{|t|_{p}^{1+p}|s|_{q}^{1+q}}dtds,$
where $f_{X,Y}$ and $f_{X},f_{Y}$ are the joint and marginal characteristic functions, respectively. Additionally, $|\cdot|_{p}$ denotes the euclidean norm in $\mathbb{R}^{p}$ and we define $c_{p}:=\pi^{(1+p)/2}/\Gamma((1+p)/2)$ with $\Gamma(\cdot)$ denoting the complete gamma function.
Thus, resembling Pearson correlation, distance correlation is the nonnegative number $\mathcal{R}(X,Y)$ defined by
$\mathcal{R}^{2}(X,Y):=\begin{cases}\frac{\mathcal{V}^{2}(X,Y)}{\sqrt{\mathcal{V}^{2}(X,X)\mathcal{V}^{2}(Y,Y)}},&\mathcal{V}^{2}(X,X)\mathcal{V}^{2}(Y,Y)>0\\ 0,&\mathcal{V}^{2}(X,X)\mathcal{V}^{2}(Y,Y)=0\end{cases}$
Simple calculations show that in the binary case, distance correlation equals the absolute value of the phi coefficient, $\mathcal{R}(\mathds{1}_{A},\mathds{1}_{B})=|\phi(\mathds{1}_{A},\mathds{1}_{B})|$. This equality already holds at the empirical level (see Definition 4 in Székely et al., (2007) for an estimator). Thus, as Cramér’s V and Tschuprow’s T, distance correlation does not indicate the direction of dependence, but inherits all the other shortcomings of the phi coefficient. As a consequence, it is an improper measure.
A second measure that is very popular and directed at functional relationships has been popularized by Chatterjee, (2021). It is defined as
$\xi(X,Y):=\frac{\int\mathrm{Var}(\mathbb{E}[\mathds{1}_{\{Y\geq t\}}|X])d\mu(t)}{\int\mathrm{Var}(\mathds{1}_{\{Y\geq t\}})d\mu(t)},$
where $\mu$ is the law of $Y$, and can be symmetrized by considering $\max\{\xi(X,Y),\xi(Y,X)\}$. However, its estimator relies on randomized breaking of ties and is therefore stochastic even for a given sample. Since the binary case is the case in which the most ties occur, this can lead to a lot of variability in the estimator. In addition to that, its population version collapses to the square of the phi coefficient, thus sharing its non-attainability: $\xi(X,Y)=\phi^{2}(X,Y)$.
Lastly, the mutual information (Shannon,, 1948) can also be used as a dependence measure if a suitable normalization is applied. The most popular one has been named the uncertainty coefficient and reads as
$U(X,Y):=2\left(\frac{H(X)+H(Y)-H(X,Y)}{H(X)+H(Y)}\right),$
where
$H(X):=-\sum_{i=1}^{k}p_{i\cdot}\ln(p_{i\cdot}),\ H(Y):=-\sum_{j=1}^{l}p_{\cdot j}\ln(p_{\cdot j}),\ \ H(X,Y):=-\sum_{i=1}^{k}\sum_{j=1}^{l}p_{ij}\ln(p_{ij})$
and $p_{i\cdot},p_{\cdot j},p_{ij}$ denote the marginal and joint probabilities of the respective values that the (discrete) random variables $X$ and $Y$ take (see the setup around equation (4)). As shown in Wermuth, (2025), it attains 1 if and only if the contingency table only has one non-zero entry in each row and each column and thus suffers from the same non-attainability problem as $\phi$.
@@ Appendix B Inference Using the Fisher Transformation
for a generic dependence measure $\delta\in\{{Q},{C},\phi,\widehat{{Q}}_{n},\widehat{{C}}_{n},\widehat{\phi}_{n}\}$. This transformation maps the bounded dependence measures to an unbounded scale while leaving values in the center (approximately the interval [0.5,0.5][-0.5,0.5]) almost unchanged. On this unbounded scale, asymptotic normality is usually a much better finite sample approximation, especially when the dependence measure is close to the boundary. We therefore apply the Fisher transformation to construct more reliable tests and confidence intervals for Q{Q} (as brought up in Bonett and Price, (2007)), C{C} and $\phi$. Even though the limiting distribution is not Gaussian for C{C}, the same problem applies here (unless C=0{C}=0).
@@
Starting with the limiting distributions of the Fisher transformation as given below in Propositions B.1B.3, we obtain non-normal limiting distributions for $\widehat{{Q}}_{n}$, $\widehat{{C}}_{n}$ and $\widehat{\phi}_{n}$ by applying the inverse transformation $Z^{-1}(z)=\mathrm{tanh}(z)=(e^{2z}-1)/(e^{2z}+1)$ of (13). These usually lead to much better approximations of the sampling distributions close to the boundaries as illustrated in Figure 6. For the true Q=0.95{Q}=0.95 with $p=q=0.5$, we plot the standard (red) and Fisher transformed (blue) limiting distributions, their implied $90\%$ confidence intervals in dashed lines, and a histogram depicting 100,000100{,}000 estimates $\widehat{{Q}}_{n}$ for a sample size of $n=100$ from simulated data. We see that the finite sample histogram is much better approximated by the blue Fisher transformed limiting distribution.
@@
Figure 6: Histogram of $\widehat{{Q}}_{n}$ from 100,000100{,}000 simulated samples of size $n=100$ with a true value of Q=0.95{Q}=0.95, marginal probabilities $p=q=0.5$ and the standard asymptotically normal and the (backtransformed) Fisher transformed limiting distributions together with corresponding $90\%$ confidence intervals in dashed lines.
@@
Consequently, the Fisher transformation leads to more reliable tests and confidence intervals close to the boundaries and almost identical results away from them. To execute a test for $\mathcal{H}_{0}:\delta=d$ one just needs to Fisher transform the hypothesized value and test $\mathcal{H}_{0}:Z(\delta)=Z(d)$ using the respective asymptotic distribution of the Fisher transformation (again using Bonferroni corrections as described above in the case of C{C}).
@@
We construct confidence intervals as follows. For Q{Q} and $\phi$, we obtain a confidence interval for the Fisher transform based on its asymptotic normality and then apply the inverse Fisher transformation to these bounds. For C{C}, we modify the algorithm described in Section 6.2 by using (inverted) tests based on the asymptotic distribution of the Fisher transforms from Proposition B.2 instead of Proposition 6.5. We illustrate the improved performance of the confidence intervals based on the Fisher transform compared to the ones based on the usual limiting distributions in simulations in Appendix D. The importance of the Fisher transformation is also illustrated when comparing the confidence intervals based on it to standard confidence intervals in the introductory example, where they differ substantially as the sampling distribution is close to the upper bound 1, see Table 6 in the Appendix, and in the case study in Section 7, where they differ in those cases and are very close to each other otherwise, see Figures 13 and 15 in the Appendix.
@@
We continue by formally providing the limiting distributions of the Fisher transformations of Q{Q}, C{C} and $\phi$ introduced in Section 6.3, which can be obtained from the asymptotic distributions from Subsection 6.1 together with the delta method. For Yule’s Q{Q}, we instead derive the limit distribution directly as in this case the transformation is a simpler expression than Yule’s Q{Q} itself, namely $Z({Q})=\frac{1}{2}\log({OR})$ with the plug-in estimator
@@
Furthermore, its first row simplifies to $\frac{1-\mathsf{q}}{2(\mathsf{p}-\mathsf{r})(1-\mathsf{p}-\mathsf{q}+\mathsf{r})}$, which is nonzero by Assumption 6.1. Hence, by the delta method (van der Vaart,, 1998, Theorem 3.1) and Lemma E.2 the claim follows. ∎
@@
  • If $\sigma=0$, then $\big(\Delta^{\top}\Omega\Delta\big)^{-1/2}\sqrt{n}\big(Z(\widehat{{C}}_{n})-Z({C})\big)\stackrel{{\scriptstyle d}}{{\to}}\frac{\gamma_{C}Z\,\mathds{1}\{Z>0\}}{m^{+}}+\frac{\gamma_{C}Z\,\mathds{1}\{Z<0\}}{m^{-}}$.
@@
    • if p=1qp=1-q, then $\sqrt{n}\big(Z(\widehat{{C}}_{n})-Z({C})\big)\stackrel{{\scriptstyle d}}{{\to}}\gamma_{C}\Lambda^{-}\begin{pmatrix}V_{4}\\ V_{3}-\mathds{1}\{V_{1}>-V_{2}\}\big[V_{1}+V_{2}\big]\end{pmatrix}$.